A factory worker places 86 newly created circuits on a shelf to be checked for quality. Of these, 6 will not work correctly. Suppose that she is asked to randomly select two circuits, without replacement, from the shelf. What is the chance that both circuits she selects will be defective? Approximate the soulution to the nearest ten thousandth.
Answer by stanbon(48516)

A factory worker places 86 newly created circuits on a shelf to be checked for quality.
Of these, 6 will not work correctly.
Suppose that she is asked to randomly select two circuits, without replacement, from the shelf.
What is the chance that both circuits she selects will be defective? Approximate the soulution to the nearest ten thousandth.
---
Number of ways to select two defective from 6: 6C2 = 15
Nomber of ways to select two with no restriction: 86C2 = (86*85)/(1*2)
43*85 = 3655
---------------------
Ans: 15/3655 = 0.0041
============================
Cheers,
Stan H.